OpenStudy (anonymous):

In Arrhenius equation Arrhenuis constant 'A' and Ea are 4*10^-13 and 19.6*10^3 if the reaction is of first order at what temperature will the half life period will be 10 times of its initial value ?

OpenStudy (aaronq):

use the equation: k=A*e^(-Ea/RT) and then use the exponential growth At=Ao*e^-kt sub in k, solve for T i'm not sure what 10 times implies, maybe 10 half-lives

OpenStudy (anonymous):

@aaronq wat is At=Ao*e^-kt

OpenStudy (anonymous):

here t is temp or time ?

OpenStudy (anonymous):

t 1/2 = 0.693/k = ~0.7/k

OpenStudy (aaronq):

it's a exponential growth/decay formula, t is time, T is temperature

OpenStudy (aaronq):

you don't need to find the half-life

OpenStudy (aaronq):

btw, time would in terms of half-lives

OpenStudy (anonymous):

Confused*

OpenStudy (aaronq):

would be*

OpenStudy (anonymous):

So wat to do with "what temperature will the half life period will be 10 times"

OpenStudy (anonymous):

10 times what ?

OpenStudy (anonymous):

of its initial value ..i guess .

OpenStudy (aaronq):

okay so, you have k=A*e^(-Ea/RT) subbing it in At=Ao*e^(-(A*e^(-Ea/RT)*t) How much is left after 10 half-lives, t= 10 Ao=1 At=(1/2)^10*Ao= 0.0009765625 At/Ao= 0.0009765625 =e^(-(A*e^(-Ea/RT)*10) you have Ea, A and R, just need to isolate T now,